# Quantitative and Predictive Folding Models from Limited Single-Molecule Data Using Simulation-Based Inference

### v1

@@ Abstract
We further compared the autocorrelation function (ACF) of the measured extension, calculated separately in the folded and unfolded states. We divided all trajectories into segments assigned either to the folded or the unfolded state. For each segment, we computed the ACF to compare simulated and experimental trajectories. As shown in figure 4C, the ACFs of simulated and experimental trajectories align well in the folded well (upper panel). However, in the unfolded well, the experimental trajectories exhibit non-single exponential decay that is not captured by the simulated trajectories (lower panel). This observation is consistent with findings from constant-position measurements on the same 30R50/T4 hairpin, which identified memory effects in the unfolded well [38]. This discrepancy may arise because a Markovian diffusion model does not fully capture the complex dynamics of an unfolded DNA hairpin [39, 40], possibly because the dynamics are not 1D in the unfolded well [41, 42].
Discussion - Accurate characterization of biomolecular folding processes remains a significant challenge in biophysics. Single-molecule force spectroscopy has enabled the observation of folding and unfolding events for individual biomolecules, providing us with detailed insights into their folding mechanism. However, instrumental artifacts and measurement noise can make the analysis of fundamental properties underlying these trajectories, such as energy landscapes, very difficult. To address these challenges, we have established a framework based on simulation-based inference (SBI) to learn quantitative folding models from single-molecule measurements at constant force. Our approach combines physics-based modeling and deep learning to approximate the Bayesian posterior from simulated data. This framework enables us to infer quantitative folding models even from limited experimental data.
@@
Here, we demonstrated that our developed framework effectively analyzes single-molecule force spectroscopy data. Using a well-studied DNA hairpin measured at constant force, we successfully learned a quantitative folding model that describes the observed data. Through this approach, we inferred free energy landscapes that match the results from established deconvolution methods while requiring significantly less data and eliminating the need for control measurements of instrumental effects. Our method also provides estimates of diffusion coefficients and linker properties. The Bayesian framework offers a key advantage: it generates complete posterior distributions that go beyond point estimates, allowing us to quantify uncertainties with testable reliability, while requiring much less data.
@@
We validated our learned folding model by comparing experimental trajectories with those simulated using the best-fitting parameters. This comparison revealed that our model reproduces the experimental data in a quantitatively robust and predictive manner. Minor discrepancies between the model and the experimental data remain, which likely indicate complex non-Markovian dynamics in the experimental data. Such deviations suggest that a more complex model may be needed to capture the underlying folding dynamics fully. Two approaches could potentially address these limitations and enhance the model’s predictive capabilities. First, we could incorporate memory effects by utilizing a more complex dynamical model, which would improve predictiveness but introduce parameters that lack direct physical interpretation. Alternatively, we could use molecular dynamics simulations as the simulator $\mathcal{M}(\theta)$, allowing us to infer molecular properties from single-molecule trajectories. While this approach offers the highest fidelity, it does so at a higher computational cost.
@@
The ability to extract statistically reliable results from limited datasets is particularly valuable for studying complex biomolecular systems where extensive data collection may be impractical or impossible. By fitting a dynamical model directly to the observed data, SBI leverages time correlations within trajectories, extracting more information from the same amount of experimental data compared to methods that ignore these temporal relationships.
Our method enables the analysis of diverse molecular systems, allowing us to efficiently construct comprehensive folding models across different biomolecular systems. This approach naturally extends to other single-molecule force spectroscopy protocols, such as constant trap experiments, potentially providing a unified framework for analyzing a wide range of experimental datasets.
Code & Data Availability - The code is available at GitHub https://github.com/covinolab/SBIsmfs and is based on the SBI-toolkit [43], a PyTorch-based implementation of simulation-based inference algorithms. Data and scripts necessary to reproduce all the results presented in this paper are freely accessible at the Zenodo repository https://zenodo.org/record/14961247.
@@
Acknowlegement - We thank Dr Attila Szabo for useful discussions and feedback. L.D. and R.C. acknowledge the support of Goethe University Frankfurt, the Frankfurt Institute of Advanced Studies, the LOEWE Center for Multiscale Modelling in Life Sciences of the state of Hesse, the CRC 1507: Membrane-associated Protein Assemblies, Machineries, and Supercomplexes (P09), and the International Max Planck Research School on Cellular Biophysics, as well as computational resources and support from the Center for Scientific Computing of the Goethe University and the Jülich Supercomputing Centre. The Flatiron Institute is a division of the Simons Foundation. This work was supported by Natural Sciences and Engineering Research Council of Canada (grant reference number RGPIN-2018-04673, to MTW).
@@ I Simulator
I.1 Brownian simulations on $G(q,x)$
The simulator $\mathcal{M}(\bm{\theta)}$ simulates the system according to Brownian dynamics. We used the Euler-Maruyama scheme to simulate Brownian dynamics.
$q(t+\Delta t)$ $=q(t)-\Delta q$ (4)
$\Delta q$ $=\beta\partial_{q}G(q,x)\cdot D_{q}\Delta t+\sqrt{2D_{q}\Delta t}\cdot R_{q}(t)$ (5)
$q(t+\Delta t)$ $=q(t)-\Delta q$ (6)
$\Delta x$ $=\beta\partial_{x}G(q,x)\cdot D_{x}\Delta t+\sqrt{2D_{x}\Delta t}\cdot R_{x}(t)$ (7)
Here, $D_{x}$ and $D_{q}$ are the diffusion coefficients along $q$ and $x$. $R_{q}(t)$ and $R_{x}(t)$ are uncorrelated Gaussian noise with zero mean and unit variance. $\Delta t$ is the integration time step. $M_{\textrm{int}}$ is the total number of integration steps, so the total simulated time is $T=M_{\textrm{int}}\Delta t$. Each simulated trajectory is sub-sampled to match the saving frequency of the observed trajectory $\bm{q}_{[1:N]}^{\textrm{exp}}$ by a factor of $\Delta\nu$ so that $\bm{q}_{[1:N]}=\{q(\Delta t\Delta\nu\cdot i)\}_{i=1}^{M_{\textrm{int}/\Delta\nu}}$.
II Modeling $G_{0}(x)$
The molecular free energy surface $G_{0}(x)$ is modeled using cubic spline interpolation. We used $N_{\textrm{node}}$ equally separated spline nodes at positions $x_{i}$. Each node has one adjustable parameter $G_{0}(x_{i})$ that determines its height, with more nodes enabling a finer description of the energy surface. To model the space between adjacent nodes, we used cubic spline functions of the form $C_{i}(x)=a_{i}+b_{i}x+c_{i}x^{2}+d_{i}x^{3}$. For $N_{\textrm{node}}$ nodes, we constructed $N_{\textrm{node}}-1$ spline functions, where $C_{i}(x)$ describes the free energy in the interval between $x_{i}$ and $x_{i+1}$. For a given set of $\{G_{0}(x_{i})\}_{i=1}^{N_{\textrm{node}}}$, the spline parameters ($a_{i}$, $b_{i}$, $c_{i}$, $d_{i}$) are fitted to ensure continuous first and second derivatives at the node junctions. Once these parameters are determined, we could efficiently evaluate both the spline function and its derivatives. To confine the system within the defined energy landscape, we introduced a positive energy offset at the boundary nodes. The first two and last two nodes were shifted upwards relative to their neighboring nodes. The free energy profiles are defined up to an additive constant.
III Neural Posterior Estimation
We employed Sequential Neural Posterior Estimation (SNPE)[29] to learn a surrogate model of the Bayesian posterior $p(\bm{\theta}|\bm{q}_{[1:N]})$. This model combines two neural networks: a density estimator $f_{\phi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))$ of parameters $\phi$ to estimate posterior density over model parameters $\bm{\theta}$, and an embedding network $S_{\Psi}(\bm{q}_{[1:N]})$ of parameters $\Psi$ that learned lower-dimensional summary statistics from the input. Both networks were trained jointly using the same loss function. The density estimator is a normalizing flow (NF), which consists of a series of invertible mappings between a base Gaussian distribution and the target posterior distribution. These mappings were parameterized by neural networks. In our case, we used the Neural Spline Flow[44] as the choice for the normalizing flow. The embedding network $S_{\Psi}(\bm{q}_{[1:N]})$ consisted of two parts. First, the input time series were featurized using transition matrices (see Timeseries featurization), which were then used as the input into a single-layer convolutional neural network. The joint network is trained on a simulated dataset $\mathcal{D}=\{\bm{\theta}_{i},\bm{q}_{[1:N],i})_{i=1}^{M}$ with $M$ samples, generated by drawing parameters $\bm{\theta}_{i}\sim p(\bm{\theta})$ from the prior and using them to run simulations with the forward model $\bm{q}_{[1:N],i}\sim p(\bm{q}_{[1:N]}|\bm{\theta}_{i})$. The surrogate posterior was trained by maximizing the average log-likelihood:
$\mathcal{L}(\phi,\Psi)=\sum_{i}^{M}log(f{\phi}(\bm{\theta}_{i}|S_{\Psi}(\bm{q}_{[1:N],i}))$ (8)
After training, the surrogate model approximates the true posterior: $f_{\phi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))\approx p(\bm{\theta}|\bm{q}_{[1:N]})$ In our case, we used SNPE [45], as it requires fewer simulations to approximate the posterior for a single observation. SNPE is not amortized. Therefore, a new surrogate posterior must be retrained for each new observation, offering greater simulation efficiency for individual inferences. The SNPE inference process consists of $K$ rounds, in each round $K_{\text{sim}}$ simulations are generated and added to the data set $\mathcal{D}$. Afterwards, the surrogate posterior is trained on the extended dataset. In the first round, parameters are drawn from the prior distribution $\bm{\theta}_{i}\sim p(\bm{\theta})$, and the surrogate posterior is trained using the loss function in Equation 8. In subsequent rounds, parameters are generated from the previously trained surrogate posterior $\bm{\theta}_{i}\sim f_{\phi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))$ and used to generate new simulations $\bm{q}_{[1:N],i}\sim p(\bm{q}_{[1:N]}|\bm{\theta}_{i})$. These new simulations, which should be more similar to the observed data $\bm{q}_{[1:N]}$, are added to the dataset $\mathcal{D}$. Specifically, we use the SNPE(C)[45] implementation from the SBI-toolkit[43] to correct for prior changes when training the density estimator over multiple rounds. After $K$ rounds of simulation and training, the final surrogate posterior approximates the true posterior: $f_{\phi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))\approx p(\bm{\theta}|\bm{q}_{[1:N]})$.
IV Timeseries featurization
We compressed the measured trajectories $\bm{q}_{[1:N]}$ into a medium-dimensional vector containing summary statistics. As summary statistics we used the elements of the estimated transition matrices at different lag times $\Delta\tau$. To estimate each transition matrix $T_{ik}(\Delta\tau)$, we segmented the trajectory $\bm{q}_{[1:N]}$ into $n_{\textrm{bins}}$ equally spaced bins. The matrix elements were populated by counting transitions between bins $i$ and $k$ at each lag time, followed by column-wise normalization. This procedure is repeated for multiple lag times to capture the system’s dynamics across different timescales.
V Priors
We factorized the prior $p(\bm{\theta})$ over the parameters $\bm{\theta}=\{D_{q}/D_{x},k_{l},G(x_{0}),...,G(x_{N_{\textrm{node}}})\}$ into the individual components: $p(\bm{\theta})=p(D_{q}/D_{x})\cdot p(k_{l})\cdot\prod_{i}^{N_{\textrm{node}}}p(G_{0}(x_{i}))$. The prior for each individual parameter is either modeled as a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ with mean value $\mu$ and variance $\sigma$, or a uniform distribution $\mathcal{U}(a,b)$ with lower bound $a$ and upper bound $b$. For the ratio of diffusion coefficients $D_{q}/D_{x}$ and linker stiffness $k_{l}$, we used log-uniform distributions. For the spline nodes, we used a multivariate Gaussian model in which the heights of the spline nodes are correlated. This allowed us to encode a rough shape of the free energy profile into the prior. The mean of the multivariate Gaussian $\bm{\mu}$ encodes the most likely a priori known free-energy profile. The covariance matrix of the multivariate Gaussian $\bm{\sigma}_{i,j}$ encodes how correlated adjacent nodes are, which allows us to enforce smoothness. The covariance matrix was constructed with the following formula:
$\bm{\sigma}_{i,j}=\alpha\cdot e^{\beta|i-j|}$
With $i$ and $j$ going from 0 to $N_{\textrm{node}}$ the number of adjustable nodes. The parameter $\alpha$ scales the covariance matrix and determines the overall width of the Gaussian, while the parameter $\beta$ determines the correlation between adjacent nodes.
VI Rate Estimation
We determined the transition rate $k$ by counting the total number of transitions $n_{t}$ (both folding and unfolding events) observed in the measured trajectories. To estimate $k$, we maximized the Poisson likelihood function $p(n_{t}\mid t,k)$, where $t$ represents the trajectory duration.
$p(n_{t}|t,k)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{(kt)^{n_{t}}\cdot e^{-kt}}{n_{t}!}$
The maximum likelihood estimate for the Poisson distribution yields $k_{max}=n_{t}/t$. To quantify the uncertainty in our rate estimate, we calculated the 68% confidence interval by integrating the Poisson likelihood function. We identified the $k$ values where the integrated and normalized likelihood encompasses 16% and 84% of the probability density. This is equivalent to a Bayesian error estimation with a uniform prior distribution.
VII Hyperparameters
[…]
VIII Experimental Setup

### v2

@@ Abstract
Finally, we compared the autocorrelation functions (ACF) of the measured extension in the folded and unfolded states (Fig. 4C). While the ACFs align well in the folded state (upper panel), the experimental trajectories in the unfolded state exhibit a non-single exponential decay not captured by our simulations (lower panel). This discrepancy is consistent with previously reported memory effects for the 30R50/T4 hairpin [44]. This suggests that a Markovian diffusion model may not fully describe the complex dynamics of an unfolded DNA hairpin, possibly because the dynamics are not 1D [48, 49, 43, 47].
To show that the same analysis framework can be applied to molecules with more complex folding, including multiple intermediate states and formation of tertiary contacts, we also analyzed the folding of an RNA riboswitch apatmer. These non-coding mRNA elements regulate gene expression via ligand-dependent structural transitions [46]. We analyzed a five-second-long constant force measurement $\bm{q}^{\mathrm{exp}}_{[1:N]}$ of the add riboswitch aptamer, whose folding pathway involves five states, where in this particular trajectory four out of five states are explored [37]. The riboswitch folds by sequentially forming two hairpins that then form a ligand-binding pocket organized by tertiary contacts (Fig. 5A). We used SNPE on $\bm{q}^{\mathrm{exp}}_{[1:N]}$ to infer a folding model of the riboswitch (Supplementary Material). The folding free energy landscape shows four distinct states, each corresponding to one of the folding intermediates we expected to observe (Fig. 5B). The landscape is consistent with results from previous single-molecule studies [37], identifying positions and energies of the potential wells and barriers that match previous analyses within error (Figure 5B, black). Finally, predictive checks validate the fitted model, as trajectories $\bm{q}_{[1:N]}$ simulated with $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}$ are consistent with the experimental measurement $\bm{q}_{[1:N]}^{\textrm{exp}}$. The potentials of mean force $G(q)$, computed from the simulated (blue) and experimental trajectory (red) also show excellent agreement (Fig. 5C)
Figure 5: Riboswitch folding. (A) Schematic illustration of the folding pathway observed during the constant force experiment. Beginning from the unfolded state U, two hairpins, P2 and P3, are formed sequentially and subsequently establish a tertiary contact, P1_U. (B) Reconstructed free energy profile. Best estimate $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}$ shown in red, and 68 % confidence interval shown in blue. Positions and energies of the potential wells and barriers (black) previously deduced from experiments [37], measured relative to state P2P3. (C) Comparison between the experimental trajectory (red) and a simulated trajectory (blue) generated using $\bm{\theta}^{\textrm{exp}}_{\textrm{MAP}}$. Potential of mean force computed from the experimental (red) and simulated trajectory (blue).
Characterizing the microscopic dynamics in biomolecular folding remains a significant challenge in biophysics. Single-molecule force spectroscopy provides direct insights into folding mechanics, though instrumental noise and artifacts complicate the extraction of properties like energy landscapes. To address these issues, we have developed a framework based on simulation-based inference that enables us to learn quantitative folding models from limited experimental data.
@@
Using a DNA hairpin measured at constant force, we learned a quantitative folding model that describes the observed data. This approach estimates free energy landscapes that match the results from established deconvolution methods while requiring significantly less data and eliminating the need for separate control measurements of instrumental artifacts. Our method also provides estimates of diffusion coefficients and linker properties. The Bayesian framework offers a key advantage: it generates complete posterior distributions that go beyond point estimates, allowing us to quantify uncertainties with testable reliability, while requiring much less data.
We validated the learned models by comparing experimental trajectories with those simulated using the best-fitting parameters, finding that our model reproduces the experimental data in a quantitatively robust and predictive manner. Minor discrepancies between the model and experimental data remain. Additional analysis using trajectories of different length (Fig. S5) suggests that these errors are unlikely to arise from having insufficient data–indeed, even 1-s trajectories of the hairpin allowed high-quality inference of the landscape. Instead, residual discrepancies likely arise from other possibilities, such as non-Markovian dynamics or position-dependent diffusion in the experimental data [44, 24]. Such deviations suggest the need for a more complex model to capture the underlying folding dynamics more fully. More complex models can, in principle, be incorporated into the SBI framework, but present additional challenges that remain to be resolved in future work. Despite the simplicity of the model used in our analysis, however, our approach was still successful when extended from a simple two-state hairpin to a more complex structure featuring multiple intermediates and tertiary contacts.
@@
To improve the predictive capabilities of the framework, two approaches could potentially be used. First, we could incorporate memory effects by utilizing a more complex dynamical model, which would improve predictiveness but introduce parameters that lack direct physical interpretation. Alternatively, we could use molecular dynamics simulations as the simulator $\mathcal{M}(\theta)$, allowing us to infer molecular properties from single-molecule trajectories. While this approach offers the highest fidelity, it does so at a higher computational cost. In general, while more complex dynamical models can in be integrated into the simulation-based inference framework, they introduce additional methodological and computational challenges that remain to be addressed in future work.
@@
Extracting statistically reliable results from limited datasets is important for investigating complex systems where extensive data collection is impractical. By fitting a dynamical model directly to observed data, SBI leverages temporal correlations, extracting more information than methods that ignore these relationships. Our method efficiently constructs comprehensive folding models for diverse biomolecular systems. It extends naturally to other single-molecule force spectroscopy protocols, such as constant trap experiments, offering a unified framework for various experimental datasets.
@@
Acknowledgement - We thank Dr Attila Szabo for useful discussions and feedback. L.D. and R.C. acknowledge the support of Goethe University Frankfurt, the Frankfurt Institute of Advanced Studies, the LOEWE Center for Multiscale Modelling in Life Sciences of the state of Hesse, the CRC 1507: Membrane-associated Protein Assemblies, Machineries, and Supercomplexes (P09), and the International Max Planck Research School on Cellular Biophysics, as well as computational resources and support from the Center for Scientific Computing of the Goethe University and the Jülich Supercomputing Centre. The Flatiron Institute is a division of the Simons Foundation. This work was supported by Natural Sciences and Engineering Research Council of Canada (grant reference number RGPIN-2018-04673, to MTW).
Data availability - The code is available at GitHub https://github.com/covinolab/SBIsmfs and is based on the SBI-toolkit [1], a PyTorch-based implementation of simulation-based inference algorithms. Data and scripts necessary to reproduce all the results presented in this paper are freely accessible at the Zenodo repository https://zenodo.org/record/18767956.
@@ I Simulator
Here, $G_{0}(x)$ describes the molecular free energy surface, including the constant pulling force, which is applied during constant-force single-molecule experiments. We modeled the free energy surface $G_{0}(x)$ using cubic spline interpolation. The parameter $k_{l}$ describes the stiffness of the harmonic linker, where (xq)(x-q) is the linker extension. We assume an anisotropic $(D_{x}\neq D_{q})$ and position-independent diffusion coefficients. Here, $D_{x}$ describes the intrinsic diffusion coefficient of the molecule, and $D_{q}$ describes the diffusion coefficient of the pulling apparatus. Because $D_{x}$ is degenerate with the barrier heights of $G_{0}(x)$, only the ratio of $D_{q}/D_{x}$ is identifiable. By simulating Brownian dynamics on the free energy surface $G(q,x)$, we obtain trajectories of the molecular extension $x(t)$ and the corresponding observed extension $q(t)$.
I.1 Brownian simulations on $G(q,x)$
The simulator $\mathcal{M}(\bm{\theta)}$ simulates the system according to Brownian dynamics. We used the Euler-Maruyama scheme to simulate the dynamics. {align} q(t + Δt) = q(t) - Δq
Δq = β∂_qG(q, x)⋅D_qΔt - 2D_qΔt⋅R_q(t)
x(t + Δt) = x(t) - Δx
Δx = β∂_xG(q, x)⋅D_xΔt - 2D_xΔt⋅R_x(t) Here, $D_{x}$ and $D_{q}$ are the diffusion coefficients along $q$ and $x$. $R_{q}(t)$ and $R_{x}(t)$ are uncorrelated Gaussian noise with zero mean and unit variance. $\Delta t$ is the integration time step. $M_{\textrm{int}}$ is the total number of integration steps, so the total simulated time is $T=M_{\textrm{int}}\Delta t$. Each simulated trajectory is sub-sampled to match the saving frequency of the observed trajectory $\bm{q}_{[1:N]}^{\textrm{exp}}$ by a factor of $\Delta\nu$ so that $\bm{q}_{[1:N]}=\{q(\Delta t\Delta\nu\cdot i)\}_{i=1}^{M_{\textrm{int}/\Delta\nu}}$.
II Modeling $G_{0}(x)$
The molecular free energy surface $G_{0}(x)$ is modeled using cubic spline interpolation. We used $N_{\textrm{node}}$ equally separated spline nodes at positions $x_{i}$. Each node has one adjustable parameter $G_{0}(x_{i})$ that determines its height, with more nodes enabling a finer description of the energy surface. To model the space between adjacent nodes, we used cubic spline functions of the form $C_{i}(x)=a_{i}+b_{i}x+c_{i}x^{2}+d_{i}x^{3}$. For $N_{\textrm{node}}$ nodes, we constructed $N_{\textrm{node}}-1$ spline functions, where $C_{i}(x)$ describes the free energy in the interval between $x_{i}$ and $x_{i+1}$. For a given set of $\{G_{0}(x_{i})\}_{i=1}^{N_{\textrm{node}}}$, the spline parameters $\{(a_{i}$, $b_{i}$, $c_{i}$, $d_{i})\}_{i=1}^{N_{\textrm{node}}}$ are fitted to ensure continuous first and second derivatives at the node junctions. Once these parameters are determined, we can efficiently evaluate both the spline function and its derivatives. To confine the system within the defined energy landscape, we introduced a positive energy offset at the boundary nodes. The first two and last two nodes were shifted upwards relative to their neighboring nodes. The free energy profiles are defined up to an additive constant.
III Neural Posterior Estimation
We employed Sequential Neural Posterior Estimation (SNPE) [28] to learn a surrogate model of the Bayesian posterior $p(\bm{\theta}|\bm{q}_{[1:N]})$. The surrogate posterior combines two neural networks: a density estimator $f_{\psi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))$ of parameters $\psi$ to estimate posterior density over model parameters $\bm{\theta}$, and an embedding network $S_{\Psi}(\bm{q}_{[1:N]})$ of parameters $\Psi$ that learned lower-dimensional summary statistics from the input. Both networks were trained jointly using the same loss function. For the density estimator, we used a normalizing flow (NF), which consists of a series of invertible mappings between a base Gaussian distribution and the target posterior distribution. These mappings were parameterized by neural networks. In our case, we used the Neural Spline Flow [14] and Masked Autoregressive Flows [40] as the choice for the normalizing flows. The embedding network $S_{\Psi}(\bm{q}_{[1:N]})$ consisted of two parts. First, the input time series were featurized using transition matrices (see Timeseries Featurization), which were then used as the input into a single-layer convolutional neural network. The joint network is trained on a simulated dataset $\mathcal{D}=\{\bm{\theta}_{i},\bm{q}_{[1:N],i})_{i=1}^{M}$ with $M$ samples, generated by drawing parameters $\bm{\theta}_{i}\sim p(\bm{\theta})$ from the prior and using them to run simulations with the forward model $\bm{q}_{[1:N],i}\sim p(\bm{q}_{[1:N]}|\bm{\theta}_{i})$. The surrogate posterior was trained by maximizing the average log-likelihood: {align} L(ψ, Ψ) = ∑_i^M log(f_ψ(θ_i | S_Ψ(q_[1:N], i)) After training, the surrogate model approximates the true posterior, $f_{\psi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))\approx p(\bm{\theta}|\bm{q}_{[1:N]})$. In our case, we used SNPE [20], as it requires fewer simulations to approximate the posterior for a single observation. SNPE is not amortized. Therefore, a new surrogate posterior must be retrained for each new observation, offering greater simulation efficiency for individual inferences. The SNPE inference process consists of $K$ rounds, in each round $K_{\text{sim}}$ simulations are generated and added to the data set $\mathcal{D}$. Afterwards, the surrogate posterior is trained on the extended dataset. In the first round, parameters are drawn from the prior distribution $\bm{\theta}_{i}\sim p(\bm{\theta})$, and the surrogate posterior is trained using the loss function in Equation III. In subsequent rounds, parameters are generated from the previously trained surrogate posterior $\bm{\theta}_{i}\sim f_{\psi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))$ and used to generate new simulations $\bm{q}_{[1:N],i}\sim p(\bm{q}_{[1:N]}|\bm{\theta}_{i})$. These new simulations, which should be more similar to the observed data $\bm{q}_{[1:N]}$, are added to the dataset $\mathcal{D}$. Specifically, we use the SNPE(C) [20] implementation from the SBI-toolkit [1] to correct for prior changes when training the density estimator over multiple rounds. After $K$ rounds of simulation and training, the final surrogate posterior approximates the true posterior: $f_{\psi}(\bm{\theta}|S_{\Psi}(\bm{q}_{[1:N]}))\approx p(\bm{\theta}|\bm{q}_{[1:N]})$.
IV Timeseries Featurization
We compressed the measured trajectories $\bm{q}_{[1:N]}$ into a medium-dimensional vector containing summary statistics. As summary statistics, we used the elements of the estimated transition matrices $T_{ik}(\Delta\tau)$ at different lag times $\Delta\tau$. To estimate each transition matrix, we segmented the trajectory $\bm{q}_{[1:N]}$ into $n_{\textrm{bins}}$ equally spaced bins. The matrix elements were populated by counting transitions between bins $i$ and $k$ at each lag time, followed by column-wise normalization. This procedure is repeated for multiple lag times to capture the system’s dynamics across different timescales. To concatenate trajectories and perform inference on multiple segments, the transition counts from each trajectory were added together and then normalized.
V Priors
We factorized the prior $p(\bm{\theta})$ over the parameters $\bm{\theta}=\{D_{q}/D_{x},k_{l},G(x_{0}),...,G(x_{N_{\textrm{node}}})\}$ into the individual components: $p(\bm{\theta})=p(D_{q}/D_{x})\cdot p(k_{l})\cdot\prod_{i}^{N_{\textrm{node}}}p(G_{0}(x_{i}))$. The prior for each individual parameter is either modeled as a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ with mean value $\mu$ and variance $\sigma$, or a uniform distribution $\mathcal{U}(a,b)$ with lower bound $a$ and upper bound $b$. For the ratio of diffusion coefficients $D_{q}/D_{x}$ and linker stiffness $k_{l}$, we used log-uniform distributions. For the spline nodes, we used a multivariate Gaussian model in which the heights of the spline nodes are correlated. This allowed us to encode a rough shape of the free energy profile into the prior. The mean of the multivariate Gaussian $\bm{\mu}$ encodes the most likely a priori known free-energy profile. The covariance matrix of the multivariate Gaussian $\bm{\sigma}_{i,j}$ encodes how correlated adjacent nodes are, which allows us to enforce smoothness. The covariance matrix was constructed with the following formula: {align*} σ_i, j = α⋅e^β|i - j| With $i$ and $j$ going from 0 to $N_{\textrm{node}}$ the number of adjustable nodes. The parameter $\alpha$ scales the covariance matrix and determines the overall width of the Gaussian, while the parameter $\beta$ determines the correlation between adjacent nodes. In principle, prior predictive checks can be used as a diagnostic tool to assess whether the chosen priors lead to unrealistic model behavior and to identify poorly specified priors.
VI Rate Estimation
We determined the transition rate $k$ by counting the total number of transitions $n_{t}$ (both folding and unfolding events) observed in the measured trajectories. To estimate $k$, we maximized the Poisson likelihood function $p(n_{t}\mid t,k)$, where $t$ represents the trajectory duration.{align*} p(n_t|t, k) = (kt)nt⋅e-ktnt! The maximum likelihood estimate for the Poisson distribution yields $k_{\mathrm{max}}=n_{t}/t$. To quantify the uncertainty in our rate estimate, we calculated the 68% confidence interval by integrating the Poisson likelihood function. We identified the $k$ values where the integrated and normalized likelihood encompasses 16% and 84% of the probability density. This is equivalent to a Bayesian error estimation with a uniform prior distribution.
VII Hyperparameters
VII.1 30R50T4 DNA hairpin
@@ VII Hyperparameters
We verified that the inferred posteriors are robust to numerical and modeling choices by performing additional analyses that varied trajectory length, numerical time discretization, and spline-based free energy parameterization. Specifically, we compared inference results obtained from analyzing a single 1-s trajectory to those from two 2-s trajectories (see Time series featurization; Fig. S5), evaluated the impact of the numerical integration step by reducing $\Delta t$ by factors of 0.5 and 0.1 using a single 2 s trajectory (Fig. S7, and assessed sensitivity to the spline representation by varying the number of spline knots (10 and 20), and altered the height of the boundaries (Fig. S8). In all cases, all other hyperparameters were kept identical to the original analysis, and no substantial changes in the inferred posteriors were observed.
VII.2 Riboswitch
Simulations were performed with a time step of $\Delta t=0.01\ \mu s$ over a total of $M_{\textrm{int}}=5\times 10^{8}$ integration steps. Trajectories were saved every $\Delta\nu=5000$ steps to match the experimental sampling rate. The molecular diffusion coefficient was set to $D_{x}=0.116\ \mathrm{nm}^{2}/\mu s^{2}$ following the same routine as before. The free energy profile was represented using 15 spline nodes, with the first and last nodes fixed at an offset of $30\penalty 10000\ k_{\textrm{B}}T$ and the second and second-last nodes fixed at $15\penalty 10000\ k_{\textrm{B}}T$, leaving $N_{\textrm{node}}=11$ adjustable nodes. The prior on the diffusion coefficient ratio $D_{q}/D_{x}$ and the linker stiffness $k_{l}$ were chosen as log-uniform distributions over $10^{-1},10^{1}]$ and $10^{-2},10^{0}]\ k_{\textrm{B}}T/\mathrm{nm}^{2}$, respectively. The spline node values were assigned a multivariate Gaussian prior with a mean given by the average experimental PMF $G(q)$. To estimate the PMF, we first applied a moving average with a window of 50 data points to smooth the trajectory, and then binned the resulting smoothed data. This approach provided an approximate shape of the free energy profile. The covariance amplitude and correlation parameters were set to $\alpha=30$ and $\beta=0.1$. Trajectories were discretized into $n_{\textrm{bins}}=50$ equally spaced bins between 552 and 578 nm, and transition matrices were computed at lag times $\Delta\tau=1,10,100,1000,$ and $20000$ steps. For inference, we employed a masked autoregressive flow (MAF) as the density estimator with a single convolutional layer as the embedding network, using the implementation provided in the SBI toolkit [1]. The MAF consisted of six transformation stages, each parameterized by a neural network with three blocks of 64 hidden units. Training was performed with a batch size of 64 and a learning rate of $2.5\times 10^{-4}$, with early stopping triggered after 15 epochs without validation loss improvement. The posterior obtained after 10 training rounds of 3000 simulations each was selected for analysis, as it yielded the best posterior predictive agreement.
VIII Generalized Jensen–Shannon Divergence
To qualitatively measure the similarity between $N$ different distributions $\bm{{p}^{i}}$, we computed the Generalized Jensen-Shannon Divergence (GJSD), which is defined as the difference between the Shannon entropy H of the average distribution and the average of the individual entropies [38, 17]:
$D_{\text{GJS}}(\bm{p}^{1},\dots,\bm{p}^{N})=H\left(\frac{1}{N}\sum_{i=1}^{N}\bm{p}^{i}\right)-\frac{1}{N}\sum_{i=1}^{N}H(\bm{p}^{i}),$ (2)
with $H(p)=\mathbb{E}[-\log(p)]$. We calculated the GJSD for the 20 posterior distributions and for their marginals regarding $D_{q}/D_{x}$ and $k_{l}$. For the full posterior distributions, we measured 1.3, while for the marginals of $D_{q}/D_{x}$ and $k_{l}$, we obtained values of 0.45 and 0.20, respectively.
To establish a baseline for the expected scale of this overlap, we compared these results to a synthetic benchmark. For this purpose, we used the posterior estimate from one of the time series and used its MAP estimate to perform 20 independent simulations. We then computed the posterior for these 20 trajectories and determined their GJSD. For the full posterior distributions of the synthetic benchmark, we obtained a GJSD of 0.14, while the marginals (Fig. S9) for $D_{x}/D_{q}$ and $k_{l}$ yielded 0.04 and 0.03.
The GJSD was computed using a Monte Carlo estimator where each distribution provides an equal contribution, following the expression[38, 17]:
$D_{\text{GJS}}\approx\frac{1}{N}\sum_{i=1}^{N}\left[\frac{1}{S}\sum_{k=1}^{S}\log\left(\frac{\bm{p}^{i}(x_{i,k})}{\frac{1}{N}\sum_{j=1}^{N}\bm{p}^{j}(x_{i,k})}\right)\right]$ (3)
where $N$ is the number of distributions and $S$ is the number of samples $x_{i,k}$ drawn from the distribution $\bm{p}^{i}$. For the full posterior, we utilized 100,000 samples. To estimate the GJSD for the marginals, we first estimated the density using a Gaussian Kernel Density Estimate (KDE) and subsequently used this to calculate the divergence. We utilized 10,000 samples per marginal to estimate the GJSD.
IX Experimental Setup
IX.1 30R50T4 DNA hairpin